OpenStudy (anonymous):

what is the simplified form of y^2-y-6/y^2-2y-3

OpenStudy (anonymous):

3) y2-y-6/y2-4 = [(y-3)(y+2)]/[(y+2)(y-2)] = (y-3)/(y-2) but y=-2 and y=+2 are not in domain y-3/y-2 with the restriction y ≠ ±2